OpenStudy (anonymous):

You are driving to a vacation spot thAt is 1500 miles away. Including rest stops, it takes you 42 hours to get to the vacation spot. You estimate that you drove at an average speed of 50 miles per hour. How many hours were you not driving?

OpenStudy (campbell_st):

well 1500/50 = driving time not driving time = 42 - answer above

OpenStudy (anonymous):

The 42 hours with reststops is just extra information. Since you are going 1500 miles away and driving at an average of 50 miles per hour while you were driving all you would do is divide 1500 by 50 which would be 30 miles total driving (that would mean you speant 12 hours at rest stops 42-30=12)

OpenStudy (texaschic101):

Distance = Speed * Time....but since we are looking for time, it can be chaned to D/S = T D/S = T distance (D) = 1500 speed (S) = 50 now sub 1500/50 = 30 hrs so if your actually driving 30 hrs and your trip took 42 hrs......42 - 30 = 12 hrs not driving